53.8 grams of a complex, diamminebis(oxalato)cobaltate(II), were dissolved in water and its
ligand field splitting energy was measured as 55.0 kJ. What is the color of the complex?
Ar: Co = 59; N = 14; H = 1; C = 12; O = 16
1
Expert's answer
2020-06-08T15:31:20-0400

Solution.

"[Co(C2O4)2(NH3)2]^{2-}"

M(complex) = 269 g/mol

"n(complex) = \frac{m}{M}"

n(complex) = 0.2 mol

"E(J/mol) = \frac{55000}{0.2} = 275000 \frac{J}{mol}"

"E = \frac{h \times c}{\lambda} = \frac{Na \times h \times c}{\lambda}"

"\lambda = \frac{Na \times h \times c}{E}"

"\lambda = \frac{6.022*10^{23} \times 6.626*10^{-34} \times3*10^8}{275000} = 4.35 \times 10^{-7} = 435 \ nm"

Since the absorption of light by the complex occurs at a wavelength of 435 nm, an additional color that determines the color of the solution will be yellow.

Answer:

yellow.